# Reducing GPU Memory Fragmentation via Spatio-Temporal Planning for Efficient Large-Scale Model Training

### v1

@@ Abstract.
The rapid scaling of large language models (LLMs) has significantly increased GPU memory pressure, which is further aggravated by training optimization techniques such as virtual pipeline and recomputation that disrupt tensor lifespans and introduce considerable memory fragmentation. Default GPU memory allocators of popular deep learning frameworks like PyTorch use online strategies without knowledge of tensor lifespans, which can waste up to 43% of memory and cause out-of-memory errors, rendering optimization techniques ineffective or even unusable.
To address this, we introduce STWeaver, a GPU memory allocator for deep learning frameworks that reduces fragmentation by exploiting the spatial and temporal regularity in memory allocation behaviors of training workloads. STWeaver introduces a novel paradigm that combines offline planning with online allocation. The offline planning leverages spatio-temporal regularities to generate a near-optimal allocation plan, while the online allocation handles complex and dynamic models such as Mixture-of-Experts (MoE). Built as a pluggable PyTorch allocator, STWeaver reduces fragmentation ratio on average by 79.2% (up to 100%) across both dense and sparse models, with negligible overhead. This enables more efficient, high-throughput training configurations and improves performance by up to 32.5%.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; April 13-16, 2026; Edinburgh, UKisbn: 978-1-4503-XXXX-X/2018/06submissionid: 192
@@ 1. Introduction
In recent years, large-scale models, particularly large language models (LLMs) (Touvron et al., 2023; Achiam et al., 2023; Jiang et al., 2023; Team et al., 2023, 2024; Bai et al., 2023; Liu et al., 2024b; Zeng et al., 2022), have demonstrated extraordinary performance in language comprehension, problem reasoning, code generation, etc. The scaling law (Kaplan et al., 2020) dictates that such powerful capabilities stem from the models’ massive parameters and training data. As a result, nowadays even a medium-sized model such as Llama-3 (Dubey et al., 2024) with 70 billion parameters requires more than 1 TB GPU/accelerator memory for training, placing heavy demands on the scarce and expensive GPU memory resource.
Additionally, current large-scale model training often employs a combination of various optimization techniques to enhance overall training efficiency. Such optimization techniques serve to either boost training throughput (Narayanan et al., 2021b; Qi et al., 2023) or reduce the theoretical GPU memory demand of the training (Chen et al., 2016; Meng et al., 2017; Le et al., 2018; Rajbhandari et al., 2020; Korthikanti et al., 2023). For instance, the Virtual Pipeline (Narayanan et al., 2021b) partitions a conventional pipeline parallel stage into several virtual stages, thereby minimizing idle periods (i.e., pipeline bubbles) inherent in pipeline parallelism. Furthermore, memory optimization techniques such as recomputation (Chen et al., 2016), tensor offloading (Meng et al., 2017; Le et al., 2018), and ZeRO (Rajbhandari et al., 2020) trade additional computation or transmission for reduced GPU memory usage.
@@
Unfortunately, the memory allocators in current deep learning frameworks, such as PyTorch (Paszke et al., 2019), struggle to efficiently handle such complex allocation patterns, leading to severe memory fragmentation (up to 43% in typical scenarios). Consequently, the actual memory consumption during training significantly exceeds the theoretical allocation requirements. The root cause of fragmentation lies in the online best-fit allocation policy adopted by the allocator in popular deep learning frameworks (e.g., PyTorch). This policy allocates a requested tensor of a certain size to the most suitable memory slot without considering the tensor’s lifespan, which is unknown to the allocator. Unpredictable deallocations lead to a discontinuous memory space, making it difficult to fit new tensors, as illustrated in Figure 1(a). Over time, this increases fragmentation as free space becomes scattered and less reusable for larger requests.
@@
More critically, the increased GPU memory consumption caused by fragmentation can slow down model training. In large-scale training, configurations with higher throughput often require more GPU memory, as shown in Figure 1(b), where each point represents a different setup, i.e., using different optimization techniques. Fragmentation reduces the amount of available GPU memory, limiting the feasibility of high-throughput configurations. When such configurations are used, fragmentation can cause actual memory usage to far exceed theoretical estimates, leading to out-of-memory (OOM) errors. As a result, model developers are forced to revert to less efficient configurations, thus reducing training efficiency (e.g., up to 24.5%).
@@
To address these problems, we propose STWeaver, a novel GPU memory allocator for deep learning frameworks to reduce fragmentation. Our approach is based on the observation that GPU memory requests exhibit strong consistency across training iterations. Therefore, by pre-assigning memory addresses before training, we can reduce fragmentation caused by online allocation in current allocators.
@@
However, optimizing memory allocation requests ahead of training meets two challenges. First, offline allocation planning is NP-hard, known as Dynamic Storage Allocation problem (Wilson et al., 1995). In large-scale model training, the number of memory requests can exceed $10^{5}$, making direct optimization intractable. To obtain a near-optimal solution within an acceptable time, we extract spatio-temporal regularities from memory allocation during training and use them to guide a grouping-based optimization. This grouping approach decomposes the time and space characteristics of memory requests, significantly reducing the complexity of the optimization problem.
@@
Second, the recent emergence of sparse models of Mixture-of-Experts (MoE) models (Jiang et al., 2023; Liu et al., 2024b, a) introduces dynamics in memory allocation patterns compared to dense models. MoE models replace MLP layers with expert layers, and decide which experts to use for each token at runtime, which results in the dynamic nature of allocation request sizes. Consequently, we cannot rely on planning of certain address for the allocation requests. To address the challenge of dynamic request sizes, we propose a hybrid paradigm that combines offline planning with online allocation. By identifying reusable regions for dynamic requests before training and performing online allocation at runtime, STWeaver supports the dynamicity of allocation requests while maintaining a low fragmentation rate.
@@
We implement STWeaver as a pluggable memory allocator for PyTorch and evaluate it across over 48 training configurations on 3 different testbeds. These configurations combine diverse dense and sparse models, model sizes, optimization techniques, microbatch sizes, and training frameworks. STWeaver reduces fragmentation memory by an average of 79.2% (up to 100%), saving up to 56.3GB GPU memory with negligible impact on end-to-end training throughput. By reducing peak GPU memory usage, it enables efficient training configurations that would otherwise trigger Out-of-Memory errors, resulting in an up to 32.5% throughput improvement. We will open source STWeaver to support more developers’ efficient large-scale training.
@@
    We propose a memory allocation paradigm for large-scale model training that combines offline planning with online allocation. STWeaver is capable of generating a near-optimal allocation plan based on spatio-temporal regularities, while effectively accommodating the dynamicity of allocation requests at runtime.
@@
    We comprehensively evaluated STWeaver using diverse training configurations on different testbeds, demonstrating its wide applicability and effectiveness. It also enables more efficient model training.
@@ 2.1. Memory-driven Parallelism and Optimization
The evolution of distributed training parallelism has been driven by the critical need to fit increasingly large models into limited GPU memory. Early data parallelism (DP) strategies, which replicate the entire model, became infeasible for large-scale training. This led to model parallelism techniques such as tensor parallelism (TP) (Shoeybi et al., 2019), which partitions weights, and pipeline parallelism (PP) (Narayanan et al., 2019, 2021a, 2021b; Qi et al., 2023), which distributes layers, each with distinct memory trade-offs. To address the escalating memory demands from models like Mixture-of-Experts (MoE) or those with long sequences, more advanced methods emerged. These include expert parallelism (EP) (Lepikhin et al., 2020) for distributing experts, sequence parallelism (SP) (Korthikanti et al., 2023) for sharding activations, and ZeRO (Rajbhandari et al., 2020) optimizations, which partition optimizer states, gradients, and even weights. This progression clearly demonstrates that efficient GPU memory usage is the central consideration shaping the design of modern parallel training systems.
While parallelism strategies help distribute memory load across GPUs, the number of available GPUs is often limited. To make training fit within GPU memory, recomputation (Chen et al., 2016) and tensor offloading (Meng et al., 2017; Le et al., 2018; Rajbhandari et al., 2021) are commonly used to reduce GPU memory usage, at the cost of slower training. Recomputation involves recalculating activation tensors within model layers during backpropagation rather than storing them, allowing for memory savings. The tensor offloading technique temporarily shifts tensors to CPU memory and retrieves them back when needed. Unfortunately, even with careful and reasonable combinations of parallelism, recomputation, and offloading, the desired training configuration often encounters the out-of-memory (OOM) error due to less effective usage of GPU memory, thus falling back to a less efficient training configuration.
@@ 2.2. Low Memory Efficiency in LLM Training
When training large models on GPUs, operators generate tensors of varying sizes and lifespans, which must be managed in GPU memory. These allocation requests pose significant challenges for memory allocators of current deep learning frameworks (e.g., PyTorch). Lacking prior knowledge of allocation patterns, allocators typically use online allocation strategy (Paszke et al., 2019). To reduce system call overhead (e.g., cudaMalloc), they often pre-allocate large caching blocks and slice out chunks based on best fit policy (Shore, 1975). Over time, this leads to fragmented memory, i.e., unallocated regions too small or scattered to satisfy new requests, known as memory fragmentation. For clarity, we define memory efficiency as the ratio of the actual allocated tensor size to the reserved GPU memory size, which is $E=\frac{M_{a}}{M_{r}}$, where $M_{a}$ is the size of allocated memory, representing the theoretical memory required under current training configuration; $M_{r}$ is the total memory reserved by the allocator, representing the actual memory usage.
Large model training often leads to severe GPU memory fragmentation, especially when complex parallelism strategies are combined with memory optimization techniques (e.g., recomputation). Figure 2 shows the memory efficiency of GPT-2 (345M parameters) (Radford et al., 2019) trained on 8 NVIDIA A800-80G GPUs under different training configurations. The baseline uses 1F1B pipeline parallelism, achieves acceptable 90% memory efficiency with 52.1 GB of reserved memory. Using Virtual Pipeline Parallelism (VPP) (Narayanan et al., 2021b) can improve training throughput. However, the utilization of VPP increases the allocated memory to 51.8 GB, and complicates the memory activities, reducing memory efficiency to 80%, leading to 59.9 GB of reserved memory. This higher memory usage can lead to OOM errors in some training scenarios. Recomputation is often used to mitigate memory requirement; however, while it reduces allocated memory, it also drops memory efficiency around 60%, causing significant memory waste. Therefore, memory fragmentation prevents logically effective approaches from achieving the expected memory reduction. Not only GPT-2, we found that many popular large models (e.g., Llama (Touvron et al., 2023), Qwen (Bai et al., 2023)) suffer from serious memory fragmentation in training (see § 8).
@@
Low Memory Efficiency Slows Training. Low memory efficiency often prevents more efficient parallelism strategies from fitting within available GPUs, which is a common challenge in large model training. As a case, we trained Qwen2.5-14B on 16 NVIDIA H200 GPUs, requiring at least 2-way tensor parallelism ($TP=2$) to fit. We selected 2-way pipeline parallelism ($PP=2$), 4-way data parallelism ($DP=4$), enabled VPP to reduce bubble ratio for better training speed, but encountered OOM. To adapt, we tried three alternatives: (1) replacing VPP with 1F1B, but still occurs OOM, (2) enabling recomputation, and (3) increasing $TP$ from 2 to 4. We also applied STWeaver to $TP=2$, $PP=2$, and $DP=4$ with virtual pipeline enabled, which successfully fit in memory. The the alternative training configurations degrade training speed by 24.5% and 7.1% compared to STWeaver respectively, highlighting the critical role of memory efficiency in enabling high-performance parallelism strategies.
@@ 2.3. Memory Behavior Insights in LLM Training
Low memory efficiency stems from complex allocation and deallocation requests, making it difficult for online allocators to minimize fragmentation. While defragmentation techniques such as block merging (Paszke et al., 2019) and virtual memory stitching (Guo et al., 2024) can help, it is either suboptimal or introduce performance overhead (see §8.3). Fortunately, large model training presents an opportunity to address this challenge. We observe that it generates a largely predictable and periodic pattern of allocation requests in both spatial and temporal dimensions, which we term allocation regularity. Although optimization techniques such as virtual pipeline and recomputation add complexity, the overall allocation behavior remains regular. This regularity can be proactively exploited by allocators to create low-fragmentation plans in advance. Notably, we identify regularity across both spatial and temporal dimensions, as detailed below.
Spatial Regularity. Modern large models are comprised of a stack of Transformer layers or identical sub-networks (Radford et al., 2019). Consequently, the size of activation tensors generated during a training iteration exhibits significant repetition, which we call spatial regularity. As shown in the Figure 3, among over 50,000 tensor allocations with $>$512-byte size in a single training iteration of Llama2-7B, there are only 32 distinct tensor sizes. Notably, with optimizations like recomputation and virtual pipeline, the regularity still persists—around 32 different sizes for $>$512-byte tensor allocations.
@@
3. STWeaver Design Overview
Figure 5. Workflow of STWeaver.
STWeaver comprises three components (Figure 5): Allocation Profiler (§4), Plan Synthesizer (§5), and Runtime Allocator (§6). To generate an ahead-of-time GPU memory allocation plan, the initial step is to use the Allocation Profiler to capture the temporal (lifespan), spatial (size), and dynamicity information of all memory (allocation or free) requests within a training iteration. The request information is then fed to the Plan Synthesizer to generate an allocation plan. To this end, the Plan Synthesizer first groups the requests to reduce planning complexity based on their spatio-temporal regularities. For static requests with fixed allocation size and lifespan, a Static Allocation Plan that minimizes memory fragmentation is generated leveraging the grouping results. To handle dynamic requests with unpredictable allocation pattern, the Plan Synthesizer finds idle spaces (termed Dynamic Reusable Space) within the Static Allocation Plan that can be reused by dynamic requests later at runtime to further reduce fragmentation. During training, the Runtime Allocator is used to perform the actual memory allocation, which consists of a Static Allocator and a Dynamic Allocator. The Static Allocator handles static requests based on the Static Allocation Plan, while the Dynamic Allocator attempts to allocate dynamic requests within the Dynamic Reusable Space if possible. For dynamic requests that cannot be accommodated by the Dynamic Reusable Space, and any unexpected requests, the Runtime Allocator falls back to a caching allocator.

### v2

# STAlloc: Enhancing Memory Efficiency in Large-Scale Model Training with Spatio-Temporal Planning

@@ Abstract.
The rapid scaling of large language models (LLMs) has significantly increased GPU memory pressure, which is further aggravated by training optimization techniques such as virtual pipeline and recomputation that disrupt tensor lifespans and introduce considerable memory fragmentation. Such fragmentation stems from the use of online GPU memory allocators in popular deep learning frameworks like PyTorch, which disregard tensor lifespans. As a result, this inefficiency can waste as much as 43% of memory and trigger out-of-memory errors, undermining the effectiveness of optimization methods.
To address this, we introduce STAlloc, a GPU memory allocator for deep learning frameworks that reduces fragmentation by exploiting the spatial and temporal regularity in memory allocation behaviors of training workloads. STAlloc introduces a novel paradigm that combines offline planning with online allocation. The offline planning leverages spatio-temporal regularities to generate a near-optimal allocation plan, while the online allocation handles complex and dynamic models such as Mixture-of-Experts (MoE). Built as a pluggable PyTorch memory allocator, STAlloc reduces fragmentation ratio on average by 85.1% (up to 100%) across both dense and MoE models, with negligible overhead. This enables more efficient, high-throughput training configurations and improves throughput performance by up to 32.5%.
Memory Defragmentation, Large Language Model, GPU, Distributed Training
journalyear: 2026copyright: ccconference: European Conference on Computer Systems; April 27–30, 2026; Edinburgh, Scotland Ukbooktitle: European Conference on Computer Systems (EUROSYS ’26), April 27–30, 2026, Edinburgh, Scotland Ukdoi: 10.1145/3767295.3769335isbn: 979-8-4007-2212-7/26/04submissionid: 192ccs: Software and its engineering Memory managementccs: Computer systems organization Parallel architecturesccs: Computing methodologies Machine learning
@@ 1. Introduction
In recent years, large-scale models, particularly large language models (LLMs) (llama; gpt4; mistral; gemini; gemma; qwen; deepseek-v3; glm), have demonstrated extraordinary performance in language comprehension, problem reasoning, code generation, etc. The scaling law (scaling-law) dictates that such powerful capabilities stem from the models’ massive parameters and training data. As a result, nowadays even a medium-sized model such as Llama-3 (llama3) with 70 billion parameters requires more than 1 TB GPU/accelerator memory for training, placing heavy demands on the scarce and expensive GPU memory resource.
Additionally, current large-scale model training often employs a combination of various optimization techniques to enhance overall training efficiency. Such optimization techniques serve to either boost training throughput (megatron2; zero-bubble-pipeline) or reduce the theoretical GPU memory demand of the training (recompute1; offload1; offload2; zero; megatron3). For instance, the Virtual Pipeline (megatron2) partitions a conventional pipeline parallel stage into several virtual stages, thereby minimizing idle periods (i.e., pipeline bubbles) inherent in pipeline parallelism. Furthermore, memory optimization techniques such as recomputation (recompute1), tensor offloading (offload1; offload2), and ZeRO (zero) trade additional computation or transmission for reduced GPU memory usage.
@@
Unfortunately, the memory allocators in current deep learning frameworks, such as PyTorch (pytorch), struggle to efficiently handle such complex allocation patterns, leading to severe memory fragmentation (up to 43% in typical scenarios). Consequently, the actual memory consumption during training significantly exceeds the theoretical allocation requirements. The root cause of fragmentation lies in the online best-fit allocation policy adopted by the allocator in popular deep learning frameworks (e.g., PyTorch). This policy allocates a requested tensor of a certain size to the most suitable memory slot without considering the tensor’s lifespan, which is unknown to the allocator. Unpredictable deallocations lead to a discontinuous memory space, making it difficult to fit new tensors, as illustrated in Figure 1(a). Over time, this increases fragmentation as free space becomes scattered and less reusable for larger requests.
@@
More critically, the increased GPU memory consumption caused by fragmentation can slow down model training. In large-scale training, configurations with higher throughput often require more GPU memory, as shown in Figure 1(b), where each point represents a different setup, i.e., using different optimization techniques. Fragmentation reduces the amount of available GPU memory, limiting the feasibility of high-throughput configurations. When such configurations are used, fragmentation can cause actual memory usage to far exceed theoretical estimates, leading to out-of-memory (OOM) errors. As a result, model developers are forced to revert to less efficient configurations with extra computation or communication, thus reducing training throughput (e.g., up to 24.5%).
@@
To address these problems, we propose STAlloc, a novel GPU memory allocator for deep learning frameworks to reduce fragmentation. Our approach is based on the observation that GPU memory requests exhibit strong consistency across training iterations. Therefore, by pre-assigning memory addresses before training, we can reduce fragmentation caused by online allocation in current allocators.
@@
However, optimizing memory allocation requests ahead of training meets two challenges. First, offline allocation planning is NP-hard, known as Dynamic Storage Allocation problem (DSA-problem). In large-scale model training, the number of memory requests can exceed $10^{5}$, making direct optimization intractable. To obtain a near-optimal solution within an acceptable time, we extract spatio-temporal regularities from memory allocation during training and use them to guide a grouping-based optimization. This grouping approach decomposes the time and space characteristics of memory requests, significantly reducing the complexity of the optimization problem.
@@
Second, the recent emergence of sparse models of Mixture-of-Experts (MoE) models (mistral; deepseek-v3; deepseek-v2) introduces dynamics in memory allocation patterns compared to dense models. MoE models replace MLP layers with expert layers, and decide which experts to use for each token at runtime, which results in the dynamic nature of allocation request sizes. Consequently, we cannot rely on planning of certain address for the allocation requests. To address the challenge of dynamic request sizes, we propose a hybrid paradigm that combines offline planning with online allocation. By identifying reusable regions for dynamic requests before training and performing online allocation at runtime, STAlloc supports the dynamicity of allocation requests while maintaining a low fragmentation rate.
@@
We implement STAlloc as a pluggable memory allocator for PyTorch and evaluate it across over 50 training configurations on three different testbeds. These configurations combine diverse dense and MoE models, model sizes, optimization techniques, microbatch sizes, and training frameworks. STAlloc reduces fragmentation memory by an average of 85.1% (up to 100%), saving up to 56.3GB GPU memory with negligible impact on end-to-end training throughput. By reducing peak GPU memory usage, it enables efficient training configurations that would otherwise trigger Out-of-Memory errors, resulting in an up to 32.5% throughput improvement. We open source STAlloc to support more developers’ efficient large-scale training [^1] https://github.com/infinigence/STAlloc.
@@
    We propose a memory allocation paradigm for large-scale model training that combines offline planning with online allocation. STAlloc is capable of generating a near-optimal allocation plan based on spatio-temporal regularities, while effectively accommodating the dynamicity of allocation requests at runtime.
@@
    We comprehensively evaluated STAlloc using diverse training configurations on different testbeds, demonstrating its wide applicability and effectiveness. It also enables more efficient model training.
@@ 2.1. Memory-driven Parallelism and Optimization
The evolution of distributed training parallelism has been driven by the critical need to fit increasingly large models into limited GPU memory. Early data parallelism (DP) strategies, which replicate the entire model, became infeasible for large-scale training. This led to model parallelism techniques such as tensor parallelism (TP) (megatron), which partitions weights, and pipeline parallelism (PP) (pipedream-1f1b; gpipe; zero-bubble-pipeline), which distributes layers, each with distinct memory trade-offs. Optimization techniques like virtual pipeline parallelism (VPP) (megatron2) further optimize pipeline scheduling to reduce bubbles and improve throughput, though the more complex scheduling often increases memory usage. To address the escalating memory demands from models like Mixture-of-Experts (MoE) or those with long sequences, more advanced methods emerged. These include expert parallelism (EP) (gshard) for distributing experts, sequence parallelism (SP) (megatron3) for sharding activations, and ZeRO (zero) optimizations, which partition optimizer states, gradients, and even weights. This progression clearly demonstrates that efficient GPU memory usage is the central consideration shaping the design of modern parallel training systems.
While parallelism strategies help distribute memory load across GPUs, the number of available GPUs is often limited. To make training fit within GPU memory, recomputation (recompute1) and tensor offloading (offload1; offload2; offload3) are commonly used to reduce GPU memory usage, at the cost of slower training. Recomputation involves recalculating activation tensors within model layers during backpropagation rather than storing them, allowing for memory savings. The tensor offloading technique temporarily shifts tensors to CPU memory and retrieves them back when needed. Unfortunately, even with careful and reasonable combinations of parallelism, recomputation, and offloading, the desired training configuration often encounters the OOM error due to less effective usage of GPU memory, thus falling back to a less efficient training configuration.
@@ 2.2. Low Memory Efficiency in LLM Training
When training large models on GPUs, operators generate tensors of varying sizes and lifespans, which must be managed in GPU memory. These allocation requests pose significant challenges for memory allocators of current deep learning frameworks (e.g., PyTorch). Lacking prior knowledge of allocation patterns, allocators typically use online allocation strategy (pytorch). To reduce system call overhead (e.g., cudaMalloc), they often pre-allocate large caching blocks and slice out chunks based on best fit policy (best-fit-allocation). Over time, this results in memory fragmentation, where free regions become too small or scattered to satisfy new allocation requests. For clarity, we define memory efficiency ($E$) as the ratio of the actual allocated tensor size to the reserved GPU memory size, which is:
(1) $E=\frac{M_{a}}{M_{r}}$
where $M_{a}$ is the size of allocated memory, representing the theoretical memory required under current training configuration; $M_{r}$ is the total memory reserved by the allocator, representing the actual memory usage.
Large model training often leads to severe GPU memory fragmentation, especially when complex parallelism strategies are combined with memory optimization techniques (e.g., recomputation). Figure 2 shows the memory efficiency of GPT-2 (345M parameters) (openai-gpt2) trained on 8 NVIDIA A800-80G GPUs under different training configurations. The baseline uses 1F1B pipeline parallelism, achieves acceptable 90% memory efficiency with 52.1 GB of reserved memory. Using Virtual Pipeline Parallelism (VPP) (megatron2) can improve training throughput. However, the utilization of VPP increases the allocated memory to 51.8 GB, and complicates the memory activities, reducing memory efficiency to 80%, leading to 59.9 GB of reserved memory. This higher memory usage can lead to OOM errors in some training scenarios. Recomputation is often used to mitigate memory requirement; however, while it reduces allocated memory, it also drops memory efficiency around 60%, causing significant memory waste. Therefore, memory fragmentation prevents logically effective approaches from achieving the expected memory reduction. Not only GPT-2, we found that many popular large models (e.g., Llama (llama), Qwen (qwen)) suffer from serious memory fragmentation in training (see § 9).
Low Memory Efficiency Slows Training. Low memory efficiency often prevents more efficient parallelism strategies from fitting within available GPUs, which is a common challenge in large model training. As a case, we trained Qwen2.5-14B on 16 NVIDIA H200 GPUs, requiring at least 2-way tensor parallelism ($TP=2$) to fit. We selected 2-way pipeline parallelism ($PP=2$), 4-way data parallelism ($DP=4$), and enabled VPP to reduce bubble ratio for better training speed, but encountered OOM. To adapt, we tried three alternatives: (1) replacing VPP with 1F1B, but still occurs OOM, (2) enabling recomputation, and (3) increasing $TP$ from 2 to 4. The alternative training configurations degrade training speed by 24.5% and 7.1% compared to the ideal case without memory fragmentation, respectively, highlighting the critical role of memory efficiency in enabling high-performance parallelism strategies.
@@ 2.3. Memory Behavior Insights in LLM Training
Low memory efficiency stems from complex allocation and deallocation requests, making it difficult for online allocators to minimize fragmentation. While defragmentation techniques such as block merging (pytorch) and virtual memory stitching (gmlake) can help, it is either suboptimal or introduce performance overhead (see §9.3). Fortunately, large model training presents an opportunity to address this challenge. We observe that it generates a largely predictable and periodic pattern of allocation requests in both spatial and temporal dimensions, which we term allocation regularity. Although optimization techniques such as virtual pipeline and recomputation add complexity, the overall allocation behavior remains regular. This regularity can be proactively exploited by allocators to create low-fragmentation plans in advance. Notably, we identify regularity across both spatial and temporal dimensions, as detailed below.
Spatial Regularity. Modern large models are comprised of a stack of Transformer layers or identical sub-networks (openai-gpt2). Consequently, the size of activation tensors generated during a training iteration exhibits significant repetition, which we call spatial regularity. As shown in the Figure 3, among over 50,000 tensor allocations with >>512-byte size in a single training iteration of Llama2-7B, there are only 32 distinct tensor sizes. Notably, with optimizations like recomputation and virtual pipeline, the regularity still persists—around 32 different sizes for >>512-byte tensor allocations.
@@
3. STAlloc Design Overview
STAlloc comprises three components (Figure 5): Allocation Profiler (§4), Plan Synthesizer (§5), and Runtime Allocator (§6). To generate an ahead-of-time GPU memory allocation plan, the initial step is to use the Allocation Profiler to capture the temporal (lifespan), spatial (size), and dynamicity information of all memory (allocation or free) requests within a training iteration. The request information is then fed to the Plan Synthesizer to generate an allocation plan. To this end, the Plan Synthesizer first groups the requests to reduce planning complexity based on their spatio-temporal regularities. For static requests with fixed allocation size and lifespan, a Static Allocation Plan that minimizes memory fragmentation is generated leveraging the grouping results. To handle dynamic requests with unpredictable allocation pattern, the Plan Synthesizer finds idle spaces (termed Dynamic Reusable Space) within the Static Allocation Plan that can be reused by dynamic requests later at runtime to further reduce fragmentation. During training, the Runtime Allocator is used to perform the actual memory allocation, which consists of a Static Allocator and a Dynamic Allocator. The Static Allocator handles static requests based on the Static Allocation Plan, while the Dynamic Allocator attempts to allocate dynamic requests within the Dynamic Reusable Space if possible. For dynamic requests that cannot be accommodated by the Dynamic Reusable Space, and any unexpected requests, the Runtime Allocator falls back to a caching allocator.
Figure 5. Workflow of STAlloc.